Your car depreciates by 8.5% each year. If you paid $9,654 for it in 2012, how much is it worth in 2016?
Answer by Cromlix(4381)

Hi there,
100% - 8.5% = 91.5%
91.5% = 91.5/100 = 0.915
2012 - 2016 = 4 years
$9,654 x 0.915^4 = $6766.93
Hope this helps :-)